# Random-Set Neural Networks (RS-NN)

### v2

@@ C.3 Algorithm for AUROC, AUPRC
[…]
Figs. 11 and 11 plot each model’s performance, illustrating the trade-offs between True Positive Rate and False Positive Rate (ROC curve), Precision and Recall (PRC curve) for CIFAR-10 (Fig. 11) and MNIST (Fig. 11). The left two plots depict the AUROC curves, where the blue curve representing RS-NN outperforms others, indicating superior discrimination between true positive and false positive rates. Similarly, on the two right plots displaying Precision-Recall Curve (PRC), RS-NN exhibits the highest curve, emphasizing its precision and recall performance. These results showcase RS-NN’s effectiveness in distinguishing in-distribution and out-of-distribution samples.
Appendix D Additional Experimental results
D.1 Robustness to noisy data
@@
Noise (scale) 0.2 0.3 0.4 0.5 0.6 0.7 0.8
Standard CNN 93.40% 79.08% 79.15% 58.33% 40.19% 28.15% 28.59%
RS-NN 96.90% 85.36% 85.91 % 68.33 % 51.46 % 37.18 % 38.05 %
LB-BNN 98.47% 95.26 % 80.18% 61.56% 43.28% 31.41% 24.55
ENN 97.81% 90.76% 75.41% 58.99% 45.70% 36.97% 31.51
Rotation (angle) -180/-120 -120/-60 -60/0 0/60 60/120 120/180 0/360
Standard CNN 36.41% 21.86% 74.41% 80.80% 23.89% 37.53% 45.86%
RS-NN 37.84% 23.54% 78.44% 81.46% 26.31% 38.48% 47.71%
LB-BNN 37.56% 20.19% 75.12% 77.67% 23.18% 37.83% 46.07
ENN 36.40% 19.43% 71.70% 78.59% 18.70% 36.38% 44.14
@@
D.2 Robustness to adversarial attacks
We conducted experiments on adversarial attacks using the well-known Fast Gradient Sign Method (FGSM) goodfellow2014explaining . FGSM is a popular approach for generating adversarial examples by perturbing input images based on the sign of the gradient of the loss function with respect to the input. Mathematically, the perturbed image $x^{\prime}$ is computed as
@@ D.2 Robustness to adversarial attacks
In our experiments, we applied FGSM adversarial attacks on the MNIST dataset for both standard CNN and Random-Set NN (RS-NN) models. This involved perturbing images from the MNIST dataset with noise generated using FGSM based on the gradient of the loss function (namely, the cross entropy loss for CNN and the $\mathcal{L}_{B-RS}$ loss, see Sec. 3.2, (7), for RS-NN).
Table 8: Test accuracies of CNN and RS-NN models under FGSM adversarial attacks on the MNIST dataset for different epsilon values. Epsilon values range from 0 to 0.05.
Model Epsilon ($\epsilon$)
@@
Subsequently, we evaluated the models’ predictions on these perturbed images and computed their test accuracies. Tab. 1 in the anonymous PDF linked to this response presents the test accuracies of both CNN and RS-NN models under FGSM adversarial attacks on the MNIST dataset for different values of $\epsilon$. As can be observed there, RS-NN demonstrates significantly higher robustness to these adversarial attacks compared to CNN, showcasing its superior resilience to adversarial perturbations. For additional information, Fig. 1 in the PDF shows examples of perturbed images generated using FSGM adversarial attacks applied to the MNIST dataset for various epsilon values ranging from 0 to 0.05.
D.3 Overconfidence in CNNs
Tab 9 shows an noisy example of a rotated MNIST image with a true label ‘$8$’. The standard CNN makes an incorrect prediction (‘$6$’) with $99.95\%$ confidence and a low entropy of $0.0061$, showcasing a noTab limitation in relying solely on confidence scores and entropy of predicted probabilities. RS-NN provides a correct prediction (class ‘$8$’) with a confidence of $49.7\%$ and pignistic entropy of $2.1626$ and a credal set width of $0.482$. Crucially, the RS-NN’s prediction process considers more than just the predicted class, taking into account both the confidence of the pignistic probability and the entropy. In this example, despite predicting the correct class (‘8’), the RS-NN demonstrates a low confidence (49.7%) and high entropy (2.1626). This holistic approach provides a more nuanced and reliable understanding of the model’s uncertainty,
True Label = "8" CNN Predictions RS-NN Predictions
Class 6 0.9995
Class 5 0.0002
Class 8 0.0001
Class 0 1.4e-05
Class 4 1.1e-06
Belief values
{’6’, ’4’, ’8’} 0.7853
{’6’, ’8’, ’1’} 0.7150
{’6’, ’8’} 0.6357
{’9’, ’8’} 0.5529
{’8’, ’3’} 0.4092
Mass values
{’6’, ’8’} 0.2079
{’9’, ’8’, ’1’} 0.1999
{’8’} 0.1959
{’8’, ’3’} 0.0961
{’6’} 0.05147
Pignistic Probability
8 0.4978
6 0.2294
9 0.1002
3 0.0502
4 0.0459
Entropy 0.0061
Entropy 2.1626 Credal set width 0.582
D.4 Budgeting of sets
The budgeting procedure can also be modified to accommodate multiple data sources or a continuous stream of data. t-SNE can be replaced by an autoencoder (this will need to be trained first) or PCA to add continual inference capabilities in the dimensionality reduction step. Similarly, Gaussian-based dynamic probabilistic clustering (GDPC) diaz2018clustering can be used instead of GMM clustering to handle a continuous stream of data. GDPC works by first initialising a GMM and then updating its parameters as it sees new data. In alternative, kulis2011revisiting have proposed an interesting approach to learn a GMM from multiple sources of data by maintaining a local and a global mixture model. Both these approaches can be plugged into our budgeting framework with little modifications to add continual learning capabilities to it. The computation of cluster overlaps remains the same, so overlapping scores and the resultant focal sets will need to be updated as the clusters evolve. However, a cluster tracking strategy barbara2001tracking can be employed so that the overlap assessment step is only re-done when a sufficient drift has been detected in the clusters. All the proposed changes are efficient enough to not have a significant effect on the overall time of budgeting.
A 2D visualization of the ellipses formed by calculating the principle axes and their lengths using eigenvectors and eigenvalues obtained from GMMs, for each class $c$ spruyt2014draw is shown in Fig. 16.
D.5 Uncertainty estimation for RS-NN
D.5.1 Entropy vs Credal Set Width
Figs. 17(a) and 17(b) depict the relationship between the entropy of RS-NN pignistic prediction and the associated confidence level (Fig. 17(a)), and between credal set width and confidence (Fig. 17(b)) for the CIFAR-10 (left) and SVHN/Intel Image (right) datasets. The distribution of iD predictions (left) for both tests show a concentration at the top left, indicating high confidence and low entropy or credal set width. Conversely, OoD predictions (middle, right) exhibit a more dispersed pattern.
Entropy, reflecting prediction uncertainty, is quite correlated with confidence, in both iD and OoD tests. In contrast, as it considers the entire set of plausible outcomes within a belief function rather than a single prediction, credal set width better quantifies the degree of epistemic uncertainty inherent to a prediction and, As a result, credal set width is less dependent on the concentration of predictions and is more reflective of the overall uncertainty encompassed by the model. Fig. 17(b) shows that, unlike entropy, credal with is clearly not correlated with confidence (Fig. 17(b)).
D.5.2 Qualitative results of entropy
Fig. 18 depicts entropy-based uncertainty estimates for rotated out-of-distribution (OoD) MNIST digit ‘$3$’ samples. All models accurately predict the true class at $-30^{\circ}$, $0^{\circ}$, and $30^{\circ}$. RS-NN and LB-BNN consistently exhibit low entropy in these scenarios, while ENN shows higher entropy for correct classifications at these angles. As the rotation angle increases, indicating more challenging scenarios, all models predict the wrong class. Notably, LB-BNN fails to exhibit a significant increase in entropy for these incorrect predictions. ENN performs relatively better than RS-NN at $60^{\circ}$ and $90^{\circ}$ rotations but has previously demonstrated high entropy levels even for accurate predictions at relatively minor rotations of $-30^{\circ}$ and $+30^{\circ}$ degrees. Overall, RS-NN provides a more reliable measure of uncertainty.
Figs. 20 and 20 show the entropy distributions for RS-NN on MNIST (iD) vs Fashion-MNIST (OoD)/Kuzushiji-MNIST (OoD), and CIFAR-10 (iD) vs SVHN (OoD)/Intel-Image (OoD), respectively. There is a clear iD vs OoD shift in entropy for RS-NN, as detailed in Sec. 4.4.
D.5.3 Entropy of pignistic predictions
@@
Table 10: The predicted belief, mass values, pignistic probabilities, and entropy for two CIFAR-10 predictions. The figure on the top (True Label = "horse") is a certain prediction with 99.9% confidence and a low entropy of 0.0017, whereas the figure on the bottom (True Label = "cat") is an uncertain prediction with 33.3% confidence and a higher entropy of 2.6955.
Tab 10 shows sample predictions by RS-NN: belief functions, masses, and pignistic predictions for given samples of CIFAR-10. The predicted belief, mass values, pignistic probabilities, and entropy are illustrated for two CIFAR-10 predictions. In the top figure, corresponding to the true label "horse," the model makes a highly confident prediction with 99.9% confidence and a low entropy of 0.0017. Conversely, the bottom figure, associated with the true label "cat," represents an uncertain prediction with 33.3% confidence and a higher entropy of 2.6955. It’s important to note that the second image is slightly unclear and poor in quality.
@@
D.5.4 Credal set width

### v3

@@ Appendix D Implementation Details
As mentioned in Sec. 4.2, the results for FSVI are different than what was reported in (Rudner et al., 2022). The lower performance with ResNet-50 in our experiments may be attributed to the model being optimized specifically for ResNet-18. We noticed the unexpectedly low results and adhered to their suggested optimizer, SGD, instead of Adam (used in all other models), to achieve optimal performance, as mentioned in Tab. 5. Additionally, to enhance performance, we also incorporated the context points from ImageNet and CIFAR-100 as recommended in Rudner et al. (2022). Furthermore, we observed that FSVI uses more extensive data augmentation than our standardized setup. For instance, their data augmentation includes random crops and higher magnitude random flips (0.5), while we only used random horizontal/vertical flips of magnitude 0.1, potentially contributing to the performance gap. While increasing model capacity by using ResNet-50 (25.6M parameters) over ResNet-18 (11.7M parameters) should theoretically improve performance, this is not always the case in practice. ResNet-50 requires more careful optimization and is more prone to overfitting, particularly when used with smaller datasets or suboptimal augmentation. This explains why ResNet-50 did not consistently outperform ResNet-18 in our implementation of the compared methods.
Appendix E Additional Experimental results
In this section, we will address the following questions based on our experimental findings:
  • Q1: Why are most models not able to differentiate between ImageNet and ImageNet-O in terms of entropy?
  • Q2: Why does RS-NN achieve better accuracy than CNN? What is the overconfidence problem in CNN?
  • Q3: How does RS-NN handle perturbed data (noisy, rotated)?
  • Q4: What empirical evidence supports the claim that RS-NN is more robust against adversarial attacks than traditional CNNs?
  • Q5: Why is the credal set considered a good measure of epistemic uncertainty?
  • Q6: What is the importance of the mass regularizers in the loss function $\mathcal{L}_{RS}$ (Sec. 3.2)? Why is it important to properly tune the regularization parameters $\alpha$ and $\beta$, and what impact does this have on the model’s performance?
  • Q7: What are the effects of budgeting on RS-NN? How was the number of budgeted focal sets $K$ chosen? How does budgeted RS-NN differ from standard RS-NN with all the subsets in the power set?
  • Q8: How does UMAP measure as a faster alternative to t-SNE for budgeting in RS-NN? How can the budgeting procedure be adapted for continuous data streams?
E.1 ImageNet vs ImageNet-O
In Tab 2, the clear separation between iD and OoD entropy is evident for RS-NN, yet for ImageNet vs ImageNet-O, all other models struggle to distinguish between the iD vs OoD entropy and the OoD dataset almost falls within the standard deviation of the iD samples. Also, in Tab. 3, the credal set widths for CIFAR-10 vs SVHN/Intel Image and MNIST vs F-MNIST/K-MNIST are distinct, but ImageNet vs ImageNet-O is not too different.
To better understand this, it is important to consider how ImageNet-O is generated. ImageNet-O is a dataset of adversarially filtered examples for ImageNet out-of-distribution detectors. It is created by removing all the ImageNet-1K samples from the full ImageNet-22K dataset. The remaining samples, which do not belong to ImageNet-1K classes, are classified by a trained model. The samples classified by the model as ImageNet-1K classes with high confidence becomes ImageNet-O. Hence, in general, it is difficult for models to make the iD vs OoD distinction as evident by uncertainty measures in Tab. 2.
We conducted experiments on Credal Set Width and Pignistic Entropy measures of RS-NN using a ViT-B-16 (Vision Transformer) backbone, as shown in Tab. 7. The pignistic entropy and credal set width show good distinction between iD and OoD measures, unlike the uncertainty measures from other baseline models in Tab. 2.
Table 7: Credal set width and entropy for RS-NN using ViT-B-16 model on ImageNet vs ImageNet-O datasets.
Datasets Credal Set Width Pignistic Entropy Softmax Entropy
RS-NN ImageNet (iD) ($\downarrow$) $0.1082\pm 0.1926$ $2.7401\pm 2.3678$ CNN $2.3442\pm 1.2243$
ImageNet-O (OoD) ($\uparrow$) $0.1948\pm 0.2455$ $4.4238\pm 2.5744$ $2.8619\pm 1.5884$
Table 8: OoD detection and uncertainty estimation performance for iD vs OoD dataset: CIFAR-10 vs CIFAR-100.
Dataset Model In-distribution Entropy ($\downarrow$) Credal Set Width ($\downarrow$) CIFAR-100
AUROC ($\uparrow$) AUPRC ($\uparrow$) Entropy ($\uparrow$) Credal Set Width ($\uparrow$)
CIFAR-10 RS-NN $0.0802\pm 0.297$ $0.0061\pm 0.039$ $88.01$ $84.93$ $0.5778\pm 0.729$ $0.0721\pm 0.165$
This is specifically due to the peculiar relation between ImageNet and ImageNet-O, and not a general problem with the models not being able to distinguish samples when they belong to the same image distribution.
To support our claim, we conducted experiments using CIFAR-10 (iD) vs CIFAR-100 (OoD) as shown in Tab. 8 below, and the results indicate that our model distinguishes between these two datasets well, as evident through the AUROC, AUPRC, entropy and credal set width. Therefore, the issue seems to be specific to ImageNet and ImageNet-O, likely due to the class imbalance in ImageNet. As per the general trend for ImageNet in Tab. 2, all models behave uncertainly for higher number of classes which makes the difference between iD and OoD less extreme. Although, RS-NN still performs better than other models on ImageNet.
E.2 RS-NN vs CNNs: accuracy and confidence
True Label =“8” CNN Predictions RS-NN Predictions
Class 6 0.9995
Class 5 0.0002
Class 8 0.0001
Class 0 1.4e-05
Class 4 1.1e-06
Belief values
{’6’, ’4’, ’8’} 0.7853
{’6’, ’8’, ’1’} 0.7150
{’6’, ’8’} 0.6357
{’9’, ’8’} 0.5529
{’8’, ’3’} 0.4092
Mass values
{’6’, ’8’} 0.2079
{’9’, ’8’, ’1’} 0.1999
{’8’} 0.1959
{’8’, ’3’} 0.0961
{’6’} 0.05147
Pignistic Probability
8 0.4978
6 0.2294
9 0.1002
3 0.0502
4 0.0459
Entropy 0.0061
Entropy 2.1626 Credal set width 0.582
Accuracy. RS-NN consistently achieves higher accuracy than standard CNN (ResNet50) on all datasets as shown in Tab. 1. This appears to attest that, as RS-NN uses a random-set framework which does not require prior assumptions and is more data driven, it can better adapt and capture the inherent complexity of the data (as sets), contributing to its superior performance in test accuracy across various datasets (Tab. 1). The training hyperparameters for RS-NN and CNN are the same, since the training procedure is quite similar.
Confidence. We further analyse how RS-NN overcomes the overconfidence problem in CNNs. Tab 9 shows an noisy example of a rotated MNIST image with a true label ‘$8$’. The standard CNN makes an incorrect prediction (‘$6$’) with $99.95\%$ confidence and a low entropy of $0.0061$, showcasing a noTab limitation in relying solely on confidence scores and entropy of predicted probabilities. RS-NN provides a correct prediction (class ‘$8$’) with a confidence of $49.7\%$ and pignistic entropy of $2.1626$ and a credal set width of $0.482$. Crucially, the RS-NN’s prediction process considers more than just the predicted class, taking into account both the confidence of the pignistic probability and the entropy. In this example, despite predicting the correct class (‘8’), the RS-NN demonstrates a low confidence (49.7%) and high entropy (2.1626). This holistic approach provides a more nuanced and reliable understanding of the model’s uncertainty.
E.3 Robustness to noisy and rotated data
@@
Noise (scale) 0.2 0.3 0.4 0.5 0.6 0.7 0.8
Standard CNN 93.40% 79.08% 79.15% 58.33% 40.19% 28.15% 28.59%
RS-NN 96.90% 85.36% 85.91 % 68.33 % 51.46 % 37.18 % 38.05 %
LB-BNN 98.47% 95.26 % 80.18% 61.56% 43.28% 31.41% 24.55
ENN 97.81% 90.76% 75.41% 58.99% 45.70% 36.97% 31.51
Rotation (angle) -180/-120 -120/-60 -60/0 0/60 60/120 120/180 0/360
Standard CNN 36.41% 21.86% 74.41% 80.80% 23.89% 37.53% 45.86%
RS-NN 37.84% 23.54% 78.44% 81.46% 26.31% 38.48% 47.71%
LB-BNN 37.56% 20.19% 75.12% 77.67% 23.18% 37.83% 46.07
ENN 36.40% 19.43% 71.70% 78.59% 18.70% 36.38% 44.14
E.4 Robustness to adversarial attacks
We conducted experiments on adversarial attacks using the well-known Fast Gradient Sign Method (FGSM) (Goodfellow et al., 2014). FGSM is a popular approach for generating adversarial examples by perturbing input images based on the sign of the gradient of the loss function with respect to the input. Mathematically, the perturbed image $x^{\prime}$ is computed as
@@ E.4 Robustness to adversarial attacks
In our experiments, we applied FGSM adversarial attacks on the MNIST dataset for standard CNN, Random-Set NN (RS-NN), LB-BNN (Bayesian) and ENN (Ensemble) models. This involved perturbing images from the MNIST dataset with noise generated using FGSM based on the gradient of the loss function (namely, the cross entropy loss for CNN and the $\mathcal{L}_{B-RS}$ loss, see Sec. 3.2, (7), for RS-NN).
Table 13: Test accuracies of CNN, RS-NN, LB-BNN and ENN models under FGSM adversarial attacks on MNIST and CIFAR-10 dataset for different epsilon values. Epsilon values range from 0 to 0.05.
Dataset Model Epsilon ($\epsilon$)
@@
Subsequently, we evaluated the models’ predictions on these perturbed images and computed their test accuracies. Tab. 13 presents the test accuracies of CNN, RS-NN, LB-BNN and ENN models under FGSM adversarial attacks on the MNIST dataset for different values of $\epsilon$.
For CIFAR-10, RS-NN demonstrates higher robustness to the attack compared to all other models with increased perturbations, while for MNIST, LB-BNN has higher accuracy at lower $\epsilon$ values but drops significantly for larger values. RS-NN circumvents the FGSM attack and shows consistent accuracy values for higher $\epsilon$. This is because FGSM for a given input image is dependent on the gradient of the loss, whereas RS-NN makes precise correct predictions with loss/gradient zero and is very confident about these predictions. As a result, the input image does not get perturbed and is correctly classified by RS-NN irrespective of the $\epsilon$ value.
Fig. 17 shows examples of perturbed images generated using FSGM adversarial attacks applied to the MNIST dataset for various epsilon values ranging from 0 to 0.05.
Additionally, we conducted experiments using the Projected Gradient Descent (PGD) adversarial attack (Madry et al., 2017), which is a more iterative and stronger attack compared to FGSM. PGD generates adversarial examples by iteratively perturbing the input image in the direction of the gradient of the loss function, followed by a projection step to ensure that the perturbation stays within a predefined range, typically defined by a norm ball around the original image. Mathematically, the perturbed image $x^{\prime}$ is computed as follows:
$x^{(0)}=x,\quad x^{(t+1)}=\text{Proj}_{\mathcal{B}(x,\epsilon)}\left(x^{(t)}+\alpha\cdot\text{sign}(\nabla_{x}J(\theta,x^{(t)},y))\right),$ (12)
where $x^{(t)}$ represents the image at the $t$-th iteration, $\alpha$ is the step size, $\epsilon$ is the maximum allowable perturbation (i.e., the size of the $\mathcal{B}(x,\epsilon)$ ball), and Proj denotes the projection operation that ensures the perturbed image $x^{\prime}$ stays within the $\epsilon$-ball around the original image $x$:
$\mathcal{B}(x,\epsilon)=\left\{x^{\prime}:\|x^{\prime}-x\|\leq\epsilon\right\}.$ (13)
After a number of iterations, typically denoted as $T$, the final perturbed image $x^{\prime}$ is obtained. This iterative process makes PGD a more powerful adversarial attack compared to FGSM. Tab. 14 presents the test accuracies of CNN, RS-NN, LB-BNN and ENN models under the PGD adversarial attack on MNIST and CIFAR-10 datasets for different values of $\epsilon$.
On MNIST, LB-BNN does well for $\epsilon=0.005$ and $\epsilon=0.01$, but CNN, LB-BNN and ENN suffer significant drops in accuracy at higher $\epsilon$ values. RS-NN performs better at higher $\epsilon$ values. On CIFAR-10, RS-NN significantly outperforms all models at all $\epsilon$ values, while the other models show a significant drop in performance, even at $\epsilon=0.005$. As explained earlier, RS-NN makes precise predictions with gradient zero and, therefore, is unaffected by PGD.
Table 14: Test accuracies of CNN, RS-NN, LB-BNN and ENN models under PGD adversarial attacks on MNIST and CIFAR-10 datasets with $\alpha$ = 1/255 and number of iterations = 10. Epsilon values range from 0 to 0.05
Dataset Model Epsilon ($\epsilon$)
0 0.005 0.01 0.02 0.03 0.04 0.05
Test acc. (%) MNIST CNN 99.12 98.90 55.44 14.03 8.11 6.58 6.58
RS-NN 99.71 99.13 93.39 91.06 89.25 89.07 89.02
LB-BNN 99.58 99.35 99.01 33.67 16.35 13.64 13.64
ENN 99.07 98.43 78.09 20.64 9.94 3.78 0.93
CIFAR-10 CNN 92.08 23.82 2.47 0.04 0 0 0
RS-NN 93.53 60.23 59.98 59.97 59.97 59.97 59.97
LB-BNN 89.95 6.28 0.51 0.55 0.59 0.59 0.59
ENN 91.55 33.95 4.40 0.15 0.02 0 0
E.5 Uncertainty estimation for RS-NN
In this section, we discuss the uncertainty measures of RS-NN: pignistic entropy (§E.5.1) and credal set width (§E.5.2), and their correlation with confidence scores (§E.5.3).
E.5.1 Entropy of pignistic predictions
The plots shown in Fig. 18 provide a comparative analysis of the entropy values across various machine learning models under two conditions: In-Distribution (iD) and Out-of-Distribution (OoD) datasets. Each subplot represents a different dataset — CIFAR-10, MNIST, and ImageNet — highlighting the performance of multiple models, including RS-NN, LB-BNN, FSVI, DE, ENN, and CNN.
Fig. 18 illustrates that RS-NN exhibits the best iD vs OoD entropy ratio among the evaluated baseline models. Specifically, RS-NN maintains significantly lower entropy values for iD datasets, indicating a strong ability to recognize familiar patterns while showing elevated entropy for OoD datasets. DE and CNN closely follow RS-NN but struggle to effectively differentiate entropy for the ImageNet dataset. LB-BNN shows more uncertainty regarding iD samples compared to OoD samples. FSVI demonstrates the weakest distinction between iD and OoD.
Figure 18: Entropy comparison of all models for iD and OoD datasets. The plots illustrate the performance of various models, with error bars indicating the standard deviation of entropy values.
Table 15: The predicted belief, mass values, pignistic probabilities, and entropy for two CIFAR-10 predictions. The figure on the top (True Label = “horse”) is a certain prediction with 99.9% confidence and a low entropy of 0.0017, whereas the figure on the bottom (True Label = “cat”) is an uncertain prediction with 33.3% confidence and a higher entropy of 2.6955.
@@ E.5.1 Entropy of pignistic predictions
Tab. 15 shows sample predictions by RS-NN: belief functions, masses, and pignistic predictions for given samples of CIFAR-10. The predicted belief, mass values, pignistic probabilities, and entropy are illustrated for two CIFAR-10 predictions. In the top figure, corresponding to the true label “horse,” the model makes a highly confident prediction with 99.9% confidence and a low entropy of 0.0017. Conversely, the bottom figure, associated with the true label “cat,” represents an uncertain prediction with 33.3% confidence and a higher entropy of 2.6955. It’s important to note that the second image is slightly unclear and poor in quality.
@@
Qualitative results of entropy. Fig. 20 depicts entropy-based uncertainty estimates for rotated out-of-distribution (OoD) MNIST digit ‘$3$’ samples. All models accurately predict the true class at $-30^{\circ}$, $0^{\circ}$, and $30^{\circ}$. RS-NN and LB-BNN consistently exhibit low entropy in these scenarios, while ENN shows higher entropy for correct classifications at these angles. As the rotation angle increases, indicating more challenging scenarios, all models predict the wrong class. Notably, LB-BNN fails to exhibit a significant increase in entropy for these incorrect predictions. ENN performs relatively better than RS-NN at $60^{\circ}$ and $90^{\circ}$ rotations but has previously demonstrated high entropy levels even for accurate predictions at relatively minor rotations of $-30^{\circ}$ and $+30^{\circ}$ degrees. Overall, RS-NN provides a more reliable measure of uncertainty.
Figs. 22 and 22 show the entropy distributions for RS-NN on MNIST (iD) vs Fashion-MNIST (OoD)/Kuzushiji-MNIST (OoD), and CIFAR-10 (iD) vs SVHN (OoD)/Intel-Image (OoD), respectively. There is a clear iD vs OoD shift in entropy for RS-NN, as detailed in Sec. 4.4.
The plots show that RS-NN consistently exhibits larger iD vs OoD entropy ratio for all datasets. For iD data, the entropy values are generally lower, indicating that these models exhibit a higher level of confidence when making predictions within familiar datasets. Conversely, the OoD side shows significantly higher entropy values for most models, particularly notable with the SVHN and Intel datasets, suggesting that the models struggle to generalize when faced with unfamiliar data. The error bars represent the standard deviation of the entropy values.
E.5.2 Credal set width